Two planes set off simultaneously flying towards each other,from airports 900km apart.one flies at 250km/h and the other at 350km/h. after how many hours will they pass each other?
Answer by stanbon(75887)

Two planes set off simultaneously flying towards each other,from airports 900km apart.one flies at 250km/h and the other at 350km/h. after how many hours will they pass each other?
----
Their closing speed is 250+350 = 600 kmh
---------------------------------
time = distance/rate
time = 900/600 hrs = 3/2 hrs = 1 hr 30 minutes
=====================================
Cheers,
Stan H.